OpenStudy (anonymous):

i need to find the solution set: 8x^2-6x-9=0

OpenStudy (anonymous):

Have you tried the general solution to the quadratic equation?

OpenStudy (anonymous):

I have tried it but im kind of confused because i missed 3 lessons so its kind of hard to understand

OpenStudy (anonymous):

\[x = -b \pm \sqrt{b ^{?}}\]

OpenStudy (anonymous):

\[x = (-b \pm \sqrt{b ^{2}-4ac})/(2a)\]

OpenStudy (anonymous):

Disregard the first equation. This is for ax^2 + bx + c = 0.

OpenStudy (anonymous):

The a,b,c lines up exactly with your question. Just substitute and solve.

OpenStudy (anonymous):

How far have you gotten?